OpenStudy (kmullis6):

What does the | symbol mean? Example: If f(x) = │x│ + 4, find f(−9). Checking the answer: f(−9) = │−9│ + 4 f(−9) = 9 + 4 f(−9) = 13 Does it mean do the opposite? -9 was changed to 9.. /:

Miracrown (miracrown):

absolute value so make the 9 +

Miracrown (miracrown):

|-9| = 9

Miracrown (miracrown):

f(-9) = |-9| + 4 = 9 + 4 = 13

OpenStudy (kmullis6):

So it does mean to make a positive a negative or a negative a positive?

OpenStudy (kmullis6):

In any case?

OpenStudy (paki):

you can take anyone first... it may be either a positive first or a negative... and V.V...

Miracrown (miracrown):

just makes whatever is inside the || lines positive

OpenStudy (kmullis6):

And if it is originally a positive, it will stay a positive?

Miracrown (miracrown):

|x| = sqrt ( x^2 ) you could rewrite it as the square root of the square

OpenStudy (kmullis6):

Or does that ever happen?

Miracrown (miracrown):

and yes...

Miracrown (miracrown):

If it is originally positive, it stays positive if it was negative, the absolute value function returns a positive

Miracrown (miracrown):

so whatever goes in comes out positive

mathslover (mathslover):

\(\LARGE x =\begin{cases} |x| = x & x > 0\\ |x| = 0& x = 0 \\ |x| = -x & x < 0\end{cases} \)

OpenStudy (kmullis6):

Thank you all so much!